OpenStudy (anonymous):

PLEASE HELP??? m'=-0.02m What is the half life? [hint: solve the equation then find at what time your solution gives half the inital mass] Here the time is measured in days

OpenStudy (unklerhaukus):

\[\frac{dm}{dt}=-0.02m\\ \frac{dm}{m}=-0.02 dt\\ \int\frac{dm}{m}=\int-0.02 dt\\ \ln m=-0.02t+c\\ m=e^{-0.02t+c}\\ \quad=m_0e^{-0.02t}\]